OpenStudy (anonymous):

solve algebraically for x: 4sin^2x=3tan^2-1 (answers: x=pi over 4, 3pi over 4, 5pi over 4, and 7pi over 4)

OpenStudy (anonymous):

is that tan^2 x?

OpenStudy (anonymous):

yes!

OpenStudy (anonymous):

Ok this you can do, Just juggle about with identites first multiply both sides by cos ^2 x, LHS becomes 4sin^2 x cos ^2 x=(sin 2x)^2. Did you understand how?

OpenStudy (anonymous):

Did you understand till here lets go step by step.

OpenStudy (anonymous):

shouldn't it be 2sin1/2^2

OpenStudy (anonymous):

Do you know the identity 2sinx cosx=sin2x

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

2sin2^2*

OpenStudy (anonymous):

good so you have here 2sinx cos x the whole squared which is sin 2x the whole squared.

OpenStudy (anonymous):

2sin2x^2**

OpenStudy (anonymous):

So On the LHS there is sin ^2 2x which we can write it as 1- cos ^2 2x

OpenStudy (anonymous):

shouldn't it be 2sin^2 2x??

OpenStudy (anonymous):

no do it agian 2 wont come.

OpenStudy (anonymous):

4sin^2xcos^2 = 2sin2x^2 not sin2x^2 ?

OpenStudy (anonymous):

no it is square in this one 2sinx cos x becomes sin 2x and another 2sinx cos x becomes sin 2x you need two 2's to make (sin x cosx)^2 (sin 2x)^2

OpenStudy (anonymous):

OHHH!!!! I didn't know that .. :P thanks, let's keep it going

OpenStudy (anonymous):

So yeah write sin ^2 2x as 1- cos ^2 2x. ok?

OpenStudy (anonymous):

kk

OpenStudy (anonymous):

Ok so now come to RHS 3tan ^2 x * cos^2 x - cos ^2 x

OpenStudy (anonymous):

tan ^2 x *cos ^2 x=sin^2 x

OpenStudy (anonymous):

so RHS becomes 3sin^2 x - cos^2 x

OpenStudy (anonymous):

Now do you know the identities sin ^2 x=(1-cos2x)/2 and cos ^2 x=(1+cos 2x)/2

OpenStudy (anonymous):

i don't know abhout the /2

OpenStudy (anonymous):

sin^2x = 1-cos^2x

OpenStudy (anonymous):

no no that identity is \[\sin ^{2}x=(1-\cos 2x)/2\]

OpenStudy (anonymous):

over 2 ?

OpenStudy (anonymous):

isn't it pythagorean indentities

OpenStudy (anonymous):

ITS \[\cos 2x \] not\[\cos ^{2}x\]

OpenStudy (anonymous):

i am so sorry. :S i didn't cover that identities yet

OpenStudy (anonymous):

those*

OpenStudy (anonymous):

You have to use those identities you have 1- cos ^2 2x on the LHS so write everything in the RHS in terms of cos 2x

OpenStudy (anonymous):

even if you didn't know them now you know them... so substitue for sin ^2 x and cos ^2 x by using the identities i told you and tell me what you get...

OpenStudy (anonymous):

did you get what i'm trying to tell?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Good so tell me what do you have on the RHS...

OpenStudy (anonymous):

unfortunately, i cannot work it out with these identities that i didn't learn for a test. .

OpenStudy (anonymous):

I'm sorry i don't know how to solve the problem in any other way.... Anyway if you want to continue with my method take LCM and all and simplify both sides you get cos ^2 2x=2cos 2x ths means cos 2x=0 and cos 2x=2(This case is not possible so the answer is just cos 2x=0

OpenStudy (anonymous):

it's okay I still appreciate you for what you've done. :)

OpenStudy (anonymous):

Okay here's an alterative easier solution , You can use sec ^2 x= 1+tan ^2 x

OpenStudy (anonymous):

So in the RHS write tan ^2 x as sec ^2 x -1 and simplify you get 8-4cos^2 x= 3sec^2 x. See if you understand till here....

OpenStudy (anonymous):

Any doubts in what i have done? If not let us say cos^2 x is some t then obviously sec ^2 x=1/cos^2 x=1/t so you get 8-4t=3/t. 8t-4t^2=3. 4t^2-8t+3=0 this implies t=1/2 or t=3/2 t=3/2 is not possible hence t=1/2 or cos ^2 x=1/2 whose obvious solutions are pi/4,3pi/4 and so on.

OpenStudy (anonymous):

lalaly wrong the equation when divided as by you becomes 4=3/cos^2x - cosec^2x

OpenStudy (lalaly):

oops sorry

OpenStudy (anonymous):

did you get it hannah?

OpenStudy (anonymous):

sorry i took a break . i am keep looking to figure out what you have done .

OpenStudy (anonymous):

i feel like you are one of my tutor that i had in my past years :) haha he was great. anyway ..

OpenStudy (anonymous):

What a compliment \m/

OpenStudy (anonymous):

Do you know how to solve quadratic equations?

OpenStudy (anonymous):

i learned it i will need to review it

OpenStudy (anonymous):

Anyway read what i have done properly its not too hard to understand i've just substitued tan^2 x as sec ^2 x -1